OpenStudy (anonymous):

For which system of equations is (2, 2) a solution?

OpenStudy (anonymous):

For which system of equations is (2, 2) a solution? A. –3x + 3y = 0 x + 6y = 10 B. –2x + 5y = –6 4x – 2y = 4 C. 5x – 2y = –6 3x – 4y = 2 D. 2x + 3y = 10 4x + 5y = 18